# Riemannian Time Warping: Multiple Sequence Alignment in Curved Spaces

### v1

@@ V-D Learning Movement Primitives with RTW pre-processing
Mean 0.4693 0.4612 1.4336 1.4292
DMP 0.4861 0.4705 1.4845 1.4560
GMM 0.4820 0.4726 1.4712 1.4682
GP 0.5648 0.5238 1.5949 1.5168
Considering all 26 characters of the dataset results in total 104 experiments. Table VII shows the average DTW distance and mean absolute error (MAE). Across all experiments, applying pre-processing with RTW significantly improves the learning result. Interestingly, the mean signal ${\mathbf{{\hat{y}}}}$ obtained from RTW yields better results than utilizing DMP, GMM and GP in this experiment.
V-E Real-Robot Teaching with Time Warping in $\mathrm{SE}(3)$
Finally, we apply time warping in real-robot teaching scenarios, utilizing a KUKA LBR iiwa manipulator with seven actuated joints. Taking advantage of the intuitive hand-guiding control mode [28], we record ${N}=4$ signals with varying lengths ${T_{n}}\in\{2048,\ldots,3667\}$ at $30$ Hz. Subsequently, the mean signal ${\mathbf{{\hat{y}}}}$ obtained from temporal alignment is executed employing a state-of-the-art QP-controller [29], taking hardware limits and manipulator dynamics into account.
The trajectory teaching is considered in $\mathrm{SE}(3)$, which is common practice in related works. We denote the end-effector pose ${\bm{x}}=\left[\begin{smallmatrix}{\bm{p}}\\ {\bm{q}}\end{smallmatrix}\right]\in{\left\{\mathbb{R}^{3}\times\mathbb{S}^{3}\right\}}$ consisting of position ${\bm{p}}\in\mathbb{R}^{3}$ and unit quaternion ${\bm{q}}\in\mathbb{S}^{3}$. Similar to [21], we concatenate the logarithmic and exponential map into single operations
${\mathbf{Log}^{\mathrm{SE}(3)}_{{\bm{e}}}\left({\bm{x}}\right)}=\left[\begin{matrix}{\bm{x}_{{\bm{p}}}}\\ {\mathbf{Log}^{\mathbb{S}^{3}}_{{\bm{e}}}\left({\bm{x}_{{\bm{q}}}}\right)}\end{matrix}\right]=\left[\begin{matrix}{\bm{u}_{{\bm{p}}}}\\ {\bm{u}_{{\bm{q}}}}\end{matrix}\right]={\bm{u}}\,,$ (22)
${\mathbf{Exp}^{\mathrm{SE}(3)}_{{\bm{e}}}\left({\bm{u}}\right)}=\left[\begin{matrix}{\bm{u}_{{\bm{p}}}}\\ {\mathbf{Exp}^{\mathbb{S}^{3}}_{{\bm{e}}}\left({\bm{u}_{{\bm{q}}}}\right)}\end{matrix}\right]=\left[\begin{matrix}{\bm{x}_{{\bm{p}}}}\\ {\bm{x}_{{\bm{q}}}}\end{matrix}\right]={\bm{x}}\,.$ (23)
In this experiment, we compare RTW with naive interpolation of the recorded data to the same length, as well as with TTW66footnotemark: 6 and NTW66footnotemark: 6 for $\mathrm{SE}(3)$ utilizing (22) and (23). Evaluation is done on two tasks, (i) following a constrained path, and (ii) repetitively drawing three circles, both shown in Figure 3.
V-E1 Following a constrained path
We instruct the robot to follow a constrained path with its end-effector, requiring precise control of position and orientation. Due to unintended time shifts within the recorded demonstration data, naive interpolation does not suffice and the resulting mean signal deviates from the desired path, resulting in a collision with the environment. TTW66footnotemark: 6, NTW66footnotemark: 6 and RTW succeed in following the constrained path without collision, however, the warping functions generated from TTW66footnotemark: 6 (15) highly compress the motion in certain sections while stretching others, resulting in a mean signal that becomes difficult to execute by the robot, as high velocities are required at the compressed sections.
V-E2 Repetitively drawing three circles
We demonstrate end-effector trajectories consisting of three circles on a whiteboard, which allows us to investigate the time warping performance for repetitive segments within provided signals. Again, naive interpolation is not sufficient. Even though both, TTW66footnotemark: 6 and NTW66footnotemark: 6, perform better, they do not align all parts of the trajectory properly. By only considering the distance between single points, TTW66footnotemark: 6 and NTW66footnotemark: 6 fail to match repetitive patterns correctly. This results in severe deviations from the desired motion at the start and end of the trajectory for both TTW66footnotemark: 6 and NTW66footnotemark: 6. Furthermore, TTW66footnotemark: 6 again suffers from fast velocity changes. Only RTW successfully aligns all parts of the demonstration data. Repetitive characteristics of the signal are correctly identified due to the Gaussian-weighted window loss (8) that considers entire segments. When executing the mean signal ${\mathbf{{\hat{y}}}}$ obtained from the proposed RTW approach, the robot’s end-effector successfully draws the three consecutive circles on the whiteboard, as shown in Figure 3.
VI Conclusion
@@
This letter introduces Riemannian Time Warping (RTW), a novel approach that efficiently aligns multiple signals while considering the geometric structure of the Riemannian manifold the data resides in, a feature that is essential for many robotic applications. We conducted extensive experiments on real world data and with an LBR iiwa robot, proofing that RTW consistently outperforms state-of-the-art baselines on both classification and averaging tasks. Furthermore, we showed that pre-processing motion data with RTW for training a robot skill improves the learning results. The computational complexity of RTW scales linearly with the number of signals and data points, which constitutes a significant advancement over techniques based on Dynamic Programming. For many robotics problems, the optimization converges after a few seconds when utilizing a high-grade GPU.

### v3

@@ V-D2 Repetitively drawing three circles
DMP 0.0392 0.0227 0.0825 0.0371
GMM 0.0302 0.0172 0.0639 0.0334
GP 0.0316 0.0175 0.0695 0.0338
TABLE VIII: Avgerage Reproduction Error (Repetitive Circles Task)
[…]
Method Raw Data RTW Data Raw Data RTW Data
DMP 0.0379 0.0222 0.0293 0.0244
GMM 0.0323 0.0220 0.0245 0.0239
GP 0.0329 0.0240 0.0245 0.0235
To further validate the benefits of RTW, we evaluate its impact on the learning performance across three imitation learning approaches: Dynamic Movement Primitives (DMP) [31], Gaussian Mixture Models (GMM) [6], Gaussian Processes (GP). Using a leave-one-out cross-validation approach, we train each method on three demonstrations and evaluate the reproduction quality on the fourth unseen demonstration. The procedure is repeated four times per task to ensure each demonstration is reproduced once. Table VII and Table VIII present the average reproduction errors for both position and orientation. The results demonstrate that RTW significantly improves the learning performance across all methods. This validates that the benefits from RTW extend beyond the time warping process itself to the entire learning pipeline, making it a valuable preprocessing tool for imitation learning in robotics.
VI Conclusion and Future Work
@@
This letter introduces Riemannian Time Warping (RTW), a novel approach that efficiently aligns multiple signals while considering the geometric structure of the Riemannian manifold the data resides in, a feature that is essential for many robotic applications. We conducted extensive experiments on synthetic and real world data, proving that RTW consistently outperforms state-of-the-art baselines on both classification and averaging tasks. Furthermore, we have shown that pre-processing motion data with RTW for training a robot primitive significantly improves the learning result. The computational complexity of RTW scales linearly with the number of signals and data points, which constitutes a significant advancement over Dynamic Programming techniques.
While RTW demonstrates strong performance across various tasks, there are limitations to be addressed in future work. RTW has only been evaluated with closed-form exponential and logarithmic maps, which do not exist for all Riemannian manifolds, wherefore numerical alternatives may be utilized that result in instabilities. Additionally, while RTW shows clear benefits for classical motion primitive learning, it remains unclear whether RTW provides similar advantages for recent imitation learning schemes like [32] or [33], which encode multi-modal motion strategies based on demonstration data with large spatial variances.